Cupcake123456:

Which is the solution to the inequality? 2 3/52 1/15 C:b< 3 2/15 B:b>3 2/15

Cupcake123456:

@dude

dude:

\(2 \frac35<b-\frac{8}{15}\)

dude:

Just for clarification purposes

dude:

Do you have to solve algebraically? Or can you graph?

Cupcake123456:

algebraically cuz im working with ( Addition and Subtraction Inequalities)

dude:

Okay, so we have to add 8/15 on both sides to isolate b

Cupcake123456:

do mark out b ?

Cupcake123456:

so*

dude:

\(2\frac{3}{5}+\frac{8}{15}\) Change the denominator (Do you know how to do that?)

dude:

\(2 \frac35<b\cancel{ \color{red}{-\frac{8}{15}}}\) \(\color{red}{+\frac{8}{15}}~~~~~~\cancel{\color{red}{+\frac{8}{15}}}\)

dude:

Yes cross out

Cupcake123456:

change to 15 ? denominator

Cupcake123456:

or can be changed to 15?

Cupcake123456:

cant*

dude:

Yeah you can't directly add \(2\frac{3}{5}+\frac{8}{15}\)

dude:

You have to change the denominator so they both have 15

Cupcake123456:

thats what i just said

dude:

Yes, do you know how to change it?

dude:

Didnt read sorry

Cupcake123456:

yes, dont u list multiples of 5 and 15? and the lowest is 15 right ?

dude:

Yes

Cupcake123456:

Okay so the next step

dude:

Convert the mixed number into an improper fraction \(2\frac35=>\frac{13}{5}\) Now we multiply the fraction by \(\frac33\) to get the denominator of 15

Cupcake123456:

and how do i do that ?

dude:

\(\large \frac{13}{5}\times\frac{3}{3}=\frac{?}{?}\)

Cupcake123456:

2 over 6 ?

dude:

Not quite, what did you do

dude:

13 * 3 =39 and 5*3 = 15 So, \(\frac{39}{15}\)

Cupcake123456:

39/15?

dude:

Right

Cupcake123456:

now.?

dude:

We can convert this back to a mixed number

Cupcake123456:

3 2/15 ?

Cupcake123456:

?????????

dude:

\(\frac{39}{15} =>2\frac{9}{15}\) Hmm it doesnt seem to be there, let me check my work

dude:

I likely did something wrong during adding and conversion I did calc and got 3.1333 aka \(3\frac{2}{15}\) \(3\frac{2}{15}<b\)